Using any data you can find in the ALEKS Data resource, calculate the equilibrium constant \(K\) at \(25.0^{\circ} \mathrm{C}\) for the following reaction.

$$ 6 \mathrm{Cl}_{2}(g)+2 \mathrm{Fe}_{2} \mathrm{O}_{3}(s) \rightarrow 4 \mathrm{FeCl}_{3}(s)+3 \mathrm{O}_{2}(g) $$

Round your answer to 2 significant digits.
Answer #1


6 de cg) + 2FC,03 (5) >> 4 Feds (s) + 30. (D saren= Daf products - sa reactant 4X-334 + 3x0) -(0+ 2x- 740.98) = -1336 + 1481.